I would like help generating kmaps with expressionscurrent if oo next br Di Toko ODO 000 - 001 010 - Lolo X ol IX 100 XL 000 00 110 000 0 ох 100 101 - 110 ir at current next ro1902 000 ool 0.0 010 010 Ol! OL) 100 100 10) 110 10 000 oo-oo-o 000 — X6 x 6 QQQ ( ooo Qe 001 olo Тоо X I olo Oil Oil 10o 100
Answer #1

0 K-map for Di, De, Jo ? ko! for Di for Dz 19.90 \QQ0 cool ll_ 10 00011 10 D2 = 4221 90+ Q29180. Di = 90+ Qe Qia for Jo for fD= 22 +229190 9-690+ @z 9 po dovko 10 1 10 Cool 11 19190 00 10 00 Ko= a1 + a2 To = P261 +2291 from 3rd truth Tash! | (3) K-ma(4 kamar for D1, D2, J.9 ks from 4th truth Table? for D2 for Di Loool it 10 1919 Pz 19190 DI = 2190T RQ,00 D2 = 9291+ 229,90